OpenStudy (anonymous):

if tan thetA=a/b then the value of a sin theta +b cos theta/a sin theta-b cos theta=

OpenStudy (anonymous):

theta=x sinx/cosx=a/b

OpenStudy (anonymous):

a *a+b *b/a*a-b*b

OpenStudy (anonymous):

a^2+b^2/a^2-b^2

OpenStudy (anonymous):

@Mandre

OpenStudy (anonymous):

@theEric

OpenStudy (souvik):

\[\frac{ a \sin \theta + b \cos \theta }{ a \sin \theta - b \cos \theta }\] divide both numerator and denominator by \(\cos \theta\) and put the value of \(\tan\theta\)

OpenStudy (anonymous):

a^2/b+1/a^2/b-1

OpenStudy (anonymous):

a^2+b/a^2-b

OpenStudy (anonymous):

same i done above

OpenStudy (souvik):

\[\frac{ \frac{ a^2 }{ b }+b }{ \frac{ a^2 }{ b }-b }\]

OpenStudy (souvik):

and then..(a^2+b^2)/(a^2-b^2)

OpenStudy (anonymous):

@souvik i have done that earlier see at the top

Parth (parthkohli):

Just because I'm a fan of fancy tricks.\[\dfrac{\sin\theta}{\cos\theta} = \dfrac{a}{b}\]\[\dfrac{a\sin\theta}{b\cos\theta} = \dfrac{a^2}{b^2}\]Applying C&D to both sides,\[\dfrac{a\sin\theta + b\cos\theta}{a\sin\theta - b\cos\theta } = \dfrac{a^2 + b^2}{a^2 - b^2}\]

ganeshie8 (ganeshie8):

clever+cute xD

OpenStudy (anonymous):

after that i have done that earlier in 3rd step

OpenStudy (anonymous):

see it required two steps.after that

OpenStudy (anonymous):

oh thats the final answer ahh

ganeshie8 (ganeshie8):

yes if u believe in C&D :)

OpenStudy (anonymous):

oh got it before itself.see i used theta =x for my convenience

OpenStudy (anonymous):

theta=x sinx/cosx=a/b

OpenStudy (anonymous):

a *a+b *b/a*a-b*b

OpenStudy (anonymous):

a^2+b^2/a^2-b^2

OpenStudy (theeric):

You guys all got it, congrats! :) Lots of individual team effort or something :)